# AIRMap — AI-Generated Radio Maps

### v1

@@ Abstract
Accurate, low-latency channel modeling is essential for real‑time wireless network simulation and digital‑twin applications. Traditional modeling methods like ray tracing are however computationally demanding and unsuited to model dynamic conditions. In this paper, we propose AIRMap, a deep‑learning framework for ultra‑fast radio‑map estimation, along with an automated pipeline for creating the largest radio‑map dataset to date. AIRMap uses a single‑input U‑Net autoencoder that processes only a 2D elevation map of terrain and building heights. Trained and evaluated on 60,000 Boston‑area samples, spanning coverage areas from 500 m to 3 km per side, AIRMap predicts path gain with under 5 dB RMSE in 4 ms per inference on an NVIDIA L40S —over $7000\times$ faster than GPU‑accelerated ray tracing based radio maps. A lightweight transfer learning calibration using just 20% of field measurements reduces the median error to approximately 10%, significantly outperforming traditional simulators, which exceed 50% error. Integration into the Colosseum emulator and the Sionna SYS platform demonstrate near-zero error in spectral efficiency and block‑error rate compared to measurement‑based channels. These findings validate AIRMap’s potential for scalable, accurate, and real‑time radio map estimation in wireless digital twins.
@@ I Introduction
In modern RAN architectures—including the Open RAN (O-RAN) framework—the threshold for “real‑time” performance varies by control‑loop function and application. Non‑real‑time RAN Intelligent Controller (RIC) applications (rApps) operate on timescales exceeding one second for long‑term optimization. Near‑real‑time functions (xApps) execute within ten to one thousand milliseconds for time‑sensitive control [9]. The most stringent tier, distributed applications (dApps), enables sub‑ten‑millisecond response times [10, 11]. Accordingly, digital‑twin systems must meet these diverse latency targets, with the strictest real‑time requirement—data acquisition and response—set below ten milliseconds for dApp‑level operations. Thus, to achieve a digital twin framework that is effective across all layers of the protocol stack, accurate and agile channel modeling is essential to precisely and swiftly characterize the radio signal propagation through a dynamic environment. This constitutes the basis of a high-fidelity digital twinning system [7].
@@
This paper presents AIRMap, a deep-learning-based framework for real-time, high-fidelity radio map estimation. AIRMap is trained on the largest site-specific dataset of its kind, automatically generated through a scalable pipeline using ray-tracing simulations and 2D elevation data. Unlike prior models requiring multiple inputs, AIRMap leverages a single-channel elevation map to produce accurate channel predictions with sub-5 dB RMSE and millisecond-level inference latency. A lightweight transfer-learning calibration procedure using sparse field measurements significantly improves accuracy, reducing median error to 10%. AIRMap is integrated into two state-of-the-art platforms–Sionna SYS and the Colosseum testbed–where it achieves near-zero error on spectral efficiency and block error rate metrics, validating its suitability for real-time digital twin applications across protocol layers.
@@
  • Large-Scale Radio Map Dataset: We develop an automated pipeline to generate the largest site-specific radio map dataset to date, comprising 60,000 Boston-area samples with diverse propagation scenarios for robust neural network training.
@@
  • Single-Input U-Net Architecture: We design a novel neural network architecture that requires only 2D elevation maps as input, achieving sub-5 dB RMSE with 4 ms inference time—over $7000\times$ faster than GPU-accelerated ray tracing.
  • Fine-tuning: We develop a novel calibration framework that leverages large-scale simulated ray-tracing data for pretraining, followed by fine-tuning with a small subset of real-world measurements. This approach effectively bridges the simulation-to-reality gap, enabling practical deployment with minimal measurement overhead while preserving model generalization and accuracy.
@@
The remainder of this paper is organized as follows. Section II evaluates conventional channel modeling techniques, focusing on ray tracing and its limitations through a real-world measurement campaign. We identify key factors—such as material properties and antenna patterns—that contribute to the gap between simulated and measured propagation data. In Section III, we present our automated pipeline for generating a large-scale dataset using Sionna RT [19], detailing trade-offs between ray-tracing fidelity and computational efficiency. We also describe the dataset structure and its suitability for deep learning. Section IV introduces AIRMap, our U-Net-based AI framework for real-time radio map estimation. We outline the model architecture, training approach, and a lightweight calibration pipeline using sparse measurements to enhance accuracy. In Section V, we demonstrate the integration of AIRMap into the Sionna SYS simulator and the Colosseum testbed, validating its accuracy and real-time performance in system-level simulations and channel emulation. Finally, Section VI concludes our contributions.
@@ II Practical Analysis of Ray-tracing-based Channel Models
For the TX side, we used an Ettus USRP X410 as the Radio Unit (RU), followed by a Minicircuits ZHL-1000-3W+ amplifier, and a Pasternack PE51OM1014 antenna. On the RX side, we employed the same antenna as used in the TX side. The RX RU was the Viavi Solutions T/Rx Software Defined Transceiver [20].
@@
Path Gain Computation from CIR. We compute the path gain directly from the measured Channel Impulse Response (CIR).
@@
Calibration techniques that adjust material properties can improve ray-tracing fidelity, but they are typically limited to small-scale indoor settings and do not extend well to large outdoor environments [22, 23, 24]. To quantify the effect of misconfigured materials and antenna patterns on path gain—especially when detailed environmental parameters are unknown—we performed a series of Sionna RT simulations. We conducted 50 different random scenarios with a maximum depth of 20 and diffraction enabled. In each run, all objects were initially assigned the ITU brick material model, and we then measured the resulting path-gain error assuming the true material was concrete. We repeated this procedure for antenna patterns by comparing TR 38.901 specifications against an isotropic radiator to isolate the impact of antenna misconfiguration. The results, summarized in Fig. 3, illustrate how incorrect material assumptions or antenna-pattern settings degrade path-gain accuracy.
@@ III-A Efficient Generation of High-Fidelity Radio Maps
Figure 6: RMSE versus runtime for varying ray‑tracing configurations in Sionna. Marker labels denote maximum path depth (reflections). Shaded regions represent 95% confidence intervals.
@@
Among various ray-tracing tools, we selected Sionna RT [19] (version 0.19.2) for its ease of automation, native Python interface, and GPU acceleration—features that streamline large-scale dataset generation. Built on top of Mitsuba 3, a differentiable rendering system powered by the just-in-time compiler Dr.Jit, Sionna RT offers a flexible and efficient framework for radio propagation modeling, making it particularly well-suited for AI-driven wireless system design.
@@
Fig. 6 shows the simulation runtime for 100 scenarios (each repeated ten times, with shaded 95% confidence intervals) across different ray-tracing configurations. As expected, enabling diffraction and increasing the maximum path depth both lead to longer runtimes. However, when diffraction is enabled, the RMSE levels off beyond a path depth of 10, indicating diminishing returns in accuracy. Fig. 7 illustrates how varying these parameters affects the resulting path-gain radio maps in a representative scenario. Based on this analysis, we identify a “sweet spot” configuration—diffraction enabled with a path depth of 10—that balances fidelity and efficiency, requiring approximately 30 seconds to generate each radio map.
@@ III-B Automated Generation of Large-Scale Datasets
For training data, we randomly select 12,000 valid TX locations within the rectangle defined by latitudes 42.2796°–42.3599° N and longitudes 71.1478°–71.0453°W, excluding points inside building footprints. Each sample comprises a pair of uniformly cropped, rasterized maps: (1) the path‑gain radio map and (2) the corresponding 2D building elevation map, with the TX centered in both. To introduce variability in coverage area, the map extent for each sample is drawn uniformly between 500 m and 3 km. This pipeline can ingest any 3D urban dataset for which detailed geometry is available, enabling seamless generalization to new locations.
To mitigate overfitting and enhance model robustness, we apply geometric data augmentation: each sample is rotated by 90°, 180°, and 270°, and flipped horizontally and vertically. This expands the dataset to 60,000 samples, offering ample diversity for training deep‑learning models. As a result, to our knowledge, this represents the largest radio‑map dataset, with effectively unlimited extensibility in both geographic scope and sample count.
@@ IV-B U-Net Architecture for Radio Map Estimation
In our approach, we define the input space as $\textbf{x}=\textbf{I}_{el}$, where $\textbf{I}_{el}$ represents a 2D elevation map derived from the scenario’s 3D model. This map encodes terrain elevation, building heights, and other physical obstructions. Although the model input has a fixed spatial dimension (e.g., $200\times 200$ pixels), we vary the spatial resolution of each sample—ranging from 2.5 m/pixel to 15 m/pixel—so that the corresponding physical area spans from 500 m to 3 km per side. This resolution-adaptive approach enables the model to learn propagation characteristics over different deployment scales
@@ IV-C Experimental Results and Analysis
Figure 11: eCDF of model performance on the test data, showing a median error of only 4.43 % across all coverage maps. Radio map values range from -150 dB to -50 dB.
For model evaluation, the entire dataset is divided using 5-fold cross-validation. Within each fold, we follow a 70-15-15 split for training, validation, and testing, respectively. To prevent data leakage, we ensure that augmented versions of the same scenario do not appear in the other sets.
As shown in Fig. 11, the median error percentage across all folds and coverage maps remains around 4.43%, demonstrating the model’s high accuracy. The path gain values in the test dataset range from -150 dB to -50 dB, encompassing a wide variety of propagation conditions. Despite this large dynamic range, the model maintains consistent performance with minimal deviations across different test cases, highlighting its robustness in estimating radio maps with exceptionally low inference error in real-time. Additionally, Fig. 10 presents qualitative results showcasing the model’s performance across various scenarios, which were randomly selected from the Boston area.
The proposed model consists of 37,601,537 parameters, reflecting its capacity to learn complex spatial patterns in radio maps. To evaluate computational efficiency, we measured the inference time over 1,000 runs on an NVIDIA L40S GPU, achieving an average of 4.2 milliseconds per run. This low-latency performance makes the model well-suited for real-time or large-scale radio map estimation in wireless systems and digital twins.
@@ IV-D Model Calibration
Given the cost and complexity of large-scale data collection, we intentionally use only a 20% subset of measurements for model calibration. This enables effective fine-tuning while preserving an unbiased portion of the data for testing. We adopt a geographically-aware train/test split that clusters points by spatial region. This approach is more appropriate for geographic data, as it enforces spatial separation between training and testing areas—thereby providing a more realistic evaluation of the model’s ability to generalize across space. To incorporate sparse measurements into training, we further modify the loss function using a weighted formulation that emphasizes measured points without overriding the knowledge learned from simulation.
@@
We perform the calibration process 100 times to account for randomness in the train-test splits, ensuring the results are not biased toward any specific subset of the data. Fig. 13 shows the performance comparison between Sionna RT, the calibrated U-Net model, and the uncalibrated model. As shown, the calibrated model consistently outperforms the others, reducing the error to around 10% and producing predictions that more closely align with real-world measurements.

### v3

# arXiv:2511.05522v3[eess.SP] 02 Mar 2026

@@ Abstract
Accurate, low-latency channel modeling is essential for real‑time wireless network simulation and digital‑twin applications. Traditional modeling methods like ray tracing are however computationally demanding and unsuited to model dynamic conditions. In this paper, we propose AIRMap, a deep‑learning framework for ultra‑fast radio‑map estimation, along with an automated pipeline for creating the largest radio‑map dataset to date. AIRMap uses a single‑input U‑Net autoencoder that processes only a 2D elevation map of terrain and building heights. Trained on 1.2M Boston-area samples and validated across four distinct urban and rural environments with varying terrain and building density, AIRMap predicts path gain with under 4 dB RMSE in 4 ms per inference on an NVIDIA L40S —over $100\times$ faster than GPU‑accelerated ray tracing based radio maps. A lightweight calibration using just 20% of field measurements reduces the median error to approximately 5%, significantly outperforming traditional simulators, which exceed 50% error. Integration into the Colosseum emulator and the Sionna SYS platform demonstrate near-zero error in spectral efficiency and block‑error rate compared to measurement‑based channels. These findings validate AIRMap’s potential for scalable, accurate, and real‑time radio map estimation in wireless digital twins.
@@ I Introduction
In modern RAN architectures—including the Open RAN (O-RAN) framework—the threshold for “real‑time” performance varies by control‑loop function and application. Non‑real‑time RAN Intelligent Controller (RIC) applications (rApps) operate on timescales exceeding one second for long‑term optimization. Near‑real‑time functions (xApps) execute within ten to one thousand milliseconds for time‑sensitive control [8]. The most stringent tier, distributed applications (dApps), enables sub‑ten‑millisecond response times [9, 28]. Accordingly, digital‑twin systems must meet these diverse latency targets, with the strictest real‑time requirement—data acquisition and response—set below ten milliseconds for dApp‑level operations. To achieve a digital twin framework that is effective across all layers of the protocol stack, accurate and agile channel modeling is essential to precisely and swiftly characterize the radio signal propagation through a dynamic environment. This constitutes the basis of a high-fidelity digital twinning system [7].
@@
This paper presents AIRMap, a deep-learning-based framework for real-time, high-fidelity radio map estimation. AIRMap is trained on the largest site-specific dataset of its kind, automatically generated through a scalable pipeline using ray-tracing simulations and 2D elevation data. Unlike prior models requiring multiple inputs, AIRMap leverages a single-channel elevation map to produce accurate channel predictions with sub-4 dB RMSE and millisecond-level inference latency. A lightweight transfer-learning calibration procedure using sparse field measurements significantly improves accuracy, reducing median error to 5%. AIRMap is integrated into two state-of-the-art platforms–Sionna SYS and the Colosseum testbed–where it achieves near-zero error on spectral efficiency and block error rate metrics, validating its suitability for real-time digital twin applications across protocol layers.
@@
  • Large-Scale Radio Map Dataset: We develop an automated pipeline to generate the largest site-specific radio map dataset to date, comprising 1.2M Boston-area samples with diverse propagation scenarios for robust neural network training.
@@
  • Single-Input U-Net Model: We design a novel neural network architecture that requires only 2D elevation maps as input, achieving sub-4 dB RMSE with 4 ms inference time, over $100\times$ faster than GPU-accelerated ray tracing.
  • Cross-Environment Generalization: We validate AIRMap across four distinct propagation environments spanning flat urban, mountainous urban, flat rural, and mountainous rural regions, demonstrating strong generalization beyond the training environment.
  • Fine-tuning: We develop a calibration framework that leverages large-scale simulated ray-tracing data for pretraining, followed by fine-tuning with a small subset of real-world measurements. This approach reduces median prediction error from over 50% to approximately 5%, effectively bridging the simulation-to-reality gap with minimal measurement overhead while preserving model generalization and accuracy.
@@
While AIRMap demonstrates strong accuracy and real-time performance, several limitations highlight important directions for future research. In its current form, the model is trained at a fixed carrier frequency of 1 GHz with isotropic antenna patterns; extending the framework to other frequency bands or directional antenna configurations may therefore require recalibration. Nevertheless, the transfer-learning strategy introduced in this work suggests that such adaptation can be efficiently achieved using limited, targeted measurement data. Moreover, AIRMap currently focuses on predicting scalar path-gain radio maps. Extending the framework to richer channel representations—such as full Channel Impulse Response (CIR), delay spread, and angular characteristics—would further enhance the fidelity of wireless digital twins. Future work will also investigate model generalization in more challenging propagation environments, including indoor–outdoor transition regions, as well as the interpretability of the learned representations to better understand how environmental features influence propagation behavior.
The remainder of this paper is organized as follows. Section II evaluates conventional channel modeling techniques, focusing on ray tracing and its limitations through a real-world measurement campaign. We identify key factors—such as material properties and antenna patterns—that contribute to the gap between simulated and measured propagation data. In Section III, we present our automated pipeline for generating a large-scale dataset using Sionna RT [20], detailing trade-offs between ray-tracing fidelity and computational efficiency. We also describe the dataset structure and its suitability for deep learning. Section IV introduces AIRMap, our U-Net-based AI framework for real-time radio map estimation. We outline the model architecture, training approach, and a lightweight calibration pipeline using sparse measurements to enhance accuracy. In Section V, we demonstrate the integration of AIRMap into the Sionna SYS simulator and the Colosseum testbed, validating its accuracy and real-time performance in system-level simulations and channel emulation. Finally, Section VI concludes our contributions. The source code for dataset generation is publicly available at https://github.com/wineslab/sionna_data_generator.
@@ II Practical Analysis of Ray-tracing-based Channel Models
For the TX side, we used an USRP X410 as the Radio Unit (RU), followed by a Minicircuits ZHL-1000-3W+ amplifier, and a Pasternack PE51OM1014 antenna. On the RX side, we employed the same antenna as used in the TX side. The RX RU was the Viavi Solutions T/Rx Software Defined Transceiver [48].
@@
Path Gain Computation from CIR. We compute the path gain directly from the measured CIR.
@@
Calibration techniques that adjust material properties can improve ray-tracing fidelity; however, such methods are typically constrained to small-scale indoor environments and do not generalize effectively to large outdoor deployments [27, 19, 26]. To systematically quantify the impact of simulation parameter misconfigurations on path-gain prediction accuracy—particularly in scenarios where detailed environmental parameters are unavailable—we conducted a comprehensive sensitivity analysis using Sionna RT. A total of 100 randomized scenarios were simulated with diffraction enabled and a maximum interaction depth of 20. Three sources of misconfiguration were evaluated independently: (i) material properties, by assigning ITU brick to all surfaces while assuming concrete as the ground-truth material; (ii) antenna radiation patterns, by comparing the 3GPP TR 38.901 specification against an isotropic radiator; and (iii) carrier frequency offset, by introducing a 70 MHz deviation from the nominal 1 GHz operating frequency. The results, summarized in Fig. 3, demonstrate the degradation in path-gain accuracy attributable to each misconfiguration source, underscoring the sensitivity of ray-tracing simulations to parameter assumptions and motivating the need for data-driven calibration approaches.
@@ III-A Efficient Generation of High-Fidelity Radio Maps
Figure 6: RMSE versus runtime for varying ray‑tracing configurations in Sionna RT (1.2.1). Marker labels denote maximum path depth (reflections).
@@
Among various ray-tracing tools, we selected Sionna RT [20] (version 1.2.1) for its ease of automation, native Python interface, and GPU acceleration—features that streamline large-scale dataset generation. Built on top of Mitsuba 3, a differentiable rendering system powered by the just-in-time compiler Dr.Jit, Sionna RT offers a flexible and efficient framework for radio propagation modeling, making it particularly well-suited for AI-driven wireless system design.
@@
Fig. 6 shows the simulation runtime for 100 scenarios (each repeated five times) across different ray-tracing configurations. As expected, enabling diffraction and increasing the maximum path depth both lead to longer runtimes. However, when diffraction is enabled, the RMSE levels off beyond a path depth of 14, indicating diminishing returns in accuracy. Fig. 7 illustrates how varying these parameters affects the resulting path-gain radio maps in a representative scenario. Based on this analysis, we identify a “sweet spot” configuration—diffraction enabled with a path depth of 14—that balances fidelity and efficiency, requiring approximately 0.265 seconds to generate each radio map.
@@ III-B Automated Generation of Large-Scale Datasets
For training data, we randomly select 240,000 valid TX locations within the rectangle defined by latitudes 42.2796°–42.3599°N and longitudes 71.1478°–71.0453°W—a region spanning approximately $100~km^{2}$ of the greater Boston area—excluding points inside building footprints. Each sample comprises a pair of uniformly cropped, rasterized maps: (1) the path-gain radio map and (2) the corresponding 2D building elevation map, with the TX centered in both. To introduce variability in coverage area, the map extent for each sample is drawn uniformly between 500 m and 3 km, corresponding to individual sample areas ranging from 0.25 km² to 9 km². This pipeline can ingest any 3D urban dataset for which detailed geometry is available, enabling seamless generalization to new locations.
To mitigate overfitting and enhance model robustness, we apply geometric data augmentation: each sample is rotated by 90°, 180°, and 270°, and flipped horizontally and vertically. Because the TX is always centered in the input image, these transformations preserve the relative spatial relationship between the transmitter and surrounding geometry—only the orientation changes. This is equivalent to rotating the physical scenario while keeping the TX-RX geometry intact, ensuring that each augmented sample represents a physically valid propagation scenario. This augmentation strategy serves two purposes: (1) it increases dataset diversity to prevent overfitting, and (2) it enforces rotational invariance, ensuring the model learns that propagation physics are independent of cardinal direction. This expands the dataset to 1.2M samples, offering ample diversity for training deep-learning models. As a result, to our knowledge, this represents the largest radio-map dataset, with effectively unlimited extensibility in both geographic scope and sample count.
@@ IV-B U-Net Architecture for Radio Map Estimation
In our approach, we define the input space as $\textbf{x}=\textbf{I}_{el}$, where $\textbf{I}_{el}$ represents a 2D elevation map derived from the scenario’s 3D model. This map encodes terrain elevation, building heights, and other physical obstructions. Although the model input has a fixed spatial dimension (e.g., $200\times 200$ pixels), we vary the spatial resolution of each sample—ranging from 2.5 m/pixel to 15 m/pixel—so that the corresponding physical area spans from 500 m to 3 km per side. This resolution-adaptive approach enables the model to learn propagation characteristics over different deployment scales.
@@ IV-C Experimental Results and Analysis
Figure 11: eCDF of model performance on the test data, showing a median error of only 3.53 % across all coverage maps. Radio map values range from -150 dB to -50 dB.
For model evaluation, we employ 5-fold cross-validation across the entire dataset. Within each fold, the data is partitioned into training, validation, and testing subsets following a 70-15-15 split ratio, respectively. To prevent data leakage, we ensure that augmented versions of the same scenario are excluded from appearing in multiple subsets. Additionally, the study area shown in Fig. 5 is divided into a $\times10$ grid, with grid cells assigned exclusively to either sets. This spatial partitioning strategy ensures that geographically proximate samples do not appear in both sets, thereby providing a more rigorous evaluation of the model’s generalization capability. As shown in Fig. 11, the median error percentage across all folds and coverage maps remains around 3.53%, demonstrating the model’s high accuracy. The path gain values in the test dataset range from -150 dB to -50 dB, encompassing a wide variety of propagation conditions. Despite this large dynamic range, the model maintains consistent performance with minimal deviations across different test cases, highlighting its robustness in estimating radio maps with exceptionally low inference error in real-time. Additionally, Fig. 10 presents qualitative results showcasing the model’s performance across various scenarios, which were randomly selected from the Boston area.
To evaluate the generalization capability of the proposed model, we conducted experiments (only testing without any calibration or retraining) on geographically and morphologically diverse regions using environmental data sourced independently from the training dataset. Specifically, building footprints were obtained from OpenStreetMap, while terrain elevation data were derived from AWS S3 public elevation tiles. Four distinct study areas were selected to represent a range of urban densities and topographical characteristics: (i) Flat coastal terrain with high-density urban development: Boston, MA; (ii) Dense urban structures in mountainous terrain: Boulder, CO; (iii) Rural mountainous region: Aspen, CO; (iv) Flat rural terrain: Lubbock, TX. The results, presented in Fig. 12, demonstrate that the model maintains sub-10% median error across most test regions, despite the use of heterogeneous data sources. As anticipated, the Boston scenario exhibits marginally higher error due to discrepancies between the OpenStreetMap-derived geometry and the BostonTwin model used during training. The Aspen scenario yields the highest error, which is attributable to the model’s limited exposure to rural mountainous environments during training. These findings confirm that AIRMap generalizes effectively across diverse geographic and morphological conditions, while also identifying domain gaps that warrant further investigation.
(a) Empirical CDF of prediction error across test regions.
(b) Satellite views illustrating the geographic and morphological diversity of the evaluated regions.
Figure 12: Generalization performance of AIRMap across diverse geographic regions. (a) eCDF of prediction error for Boston, MA (coastal urban), Boulder, CO (mountainous urban), Aspen, CO (mountainous rural), and Lubbock, TX (flat rural), evaluated using independently sourced environmental data from OpenStreetMap and AWS S3 public elevation tiles.
The proposed model comprises approximately 37.6 million trainable parameters with a memory footprint of 144.08 MB, reflecting sufficient capacity to learn complex spatial patterns inherent in radio propagation phenomena. To evaluate computational efficiency, we conducted inference benchmarks over 1,000 forward passes on an NVIDIA L40S GPU. The model achieves an average inference latency of 4.2 ms per sample, corresponding to a throughput of 238 samples/s, with a peak GPU memory utilization of 460 MB and an energy consumption of approximately 1.52 J per inference. These performance characteristics demonstrate that AIRMap is well-suited for real-time radio map estimation in wireless digital twin applications, satisfying the sub-10 ms latency requirements for dApp-level operations in O-RAN architectures while maintaining a computationally efficient footprint suitable for large-scale deployment.
@@ IV-D Model Calibration
To address simulation-to-reality discrepancies and improve real-world accuracy, we introduce a calibration pipeline that fine-tunes the pretrained model using a subset of field measurements. To ensure a realistic assessment of spatial generalization, we adopt a geographically-aware train/test split that clusters measurement points by spatial region, thereby enforcing separation between training and testing areas. This approach is particularly appropriate for geographic data, as it evaluates the model’s ability to generalize across space rather than interpolate between nearby points.
To incorporate sparse field measurements into the pretrained model, we employ an analytical calibration approach that computes an optimal affine transformation in closed form. Given the pretrained model’s predictions $\hat{P}$ and sparse measurements $P_{m}$ at $N$ measurement locations, we seek scale $\alpha$ and offset $\beta$ parameters that minimize the squared error:
$\min_{\alpha,\beta}\sum_{i=1}^{N}\left(\alpha\hat{P}_{i}+\beta-P_{m,i}\right)^{2}$ (3)
The closed-form solution is obtained via ordinary least squares:
$\alpha=\frac{\text{Cov}(\hat{P},P_{m})}{\text{Var}(\hat{P})},\quad\beta=\bar{P}_{m}-\alpha\bar{\hat{P}}$ (4)
where $\bar{\hat{P}}$ and $\bar{P}_{m}$ denote the sample means. The calibrated prediction is then $P_{\text{cal}}=\alpha\hat{P}+\beta$.
This approach offers several advantages over gradient-based fine-tuning: (i) it requires no iterative optimization, (ii) it is robust to overfitting with extremely sparse data, and (iii) it preserves the spatial patterns learned during pretraining while correcting for systematic bias between simulation and measurement domains.
@@
We perform the calibration process 100 times to account for randomness in the train-test splits, ensuring that the results are not biased toward any specific subset of the data. Fig. 14 presents a performance comparison among Sionna RT, the calibrated U-Net model, and the uncalibrated model. As shown, the calibrated model consistently outperforms the others, reducing the median error to approximately 5% and producing predictions that more closely align with real-world measurements.
The effect of varying the calibration data ratio is illustrated in Fig. 15. As shown, increasing the proportion of measurement data used for calibration yields only marginal improvements in accuracy. For instance, using 20% of the measurements achieves a median error of approximately 5%, whereas using 80% reduces the error to 3.57%—an improvement of only 1.43 percentage points. This trade-off must be considered in the context of data collection costs and availability. Furthermore, as discussed in Section V, errors of this magnitude do not significantly impact system-level performance, suggesting that 20% represents a practical and cost-effective calibration threshold.
@@
Figure 15: eCDF of the error for different proportions of measurement data used to train the calibration process.